OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) x + 6y = 1 2x + 5y = 9

OpenStudy (anonymous):

you take the soulution and times it by 3 and you have our anwer!! you r welcome

OpenStudy (anonymous):

solve for x in the first equation to get x=1-6y. substitute the x in the first equation with this to get 2(1-6y) + 5y = 9. distribute the 2 into the binomial to get the equation: 2-12y+5y=9 combine like terms: 2-7y=9 add 7y and subtract 9 from both sides to get -7=7y. divide by 7 to get -1=y